OpenStudy (anonymous):

HELPPP!!! Will give medal:) Kayla currently has an account balance of $2,134.96. She opened the account 12 years ago with a deposit of $1,327.35. If the interest compounds twice a month, what is the interest rate on the account?

OpenStudy (jdoe0001):

\(\bf A=P\left(1+\frac{r}{n}\right)^{nt} \\ \quad \\ A=\textit{current amount}\to 2,134.96\\ P=\textit{original amount deposited}\to \$1,237.35\\ r=rate\\ n=\textit{times it compounds per year, twice a month, bi-weekly}\to 24\\ t=years\to 12\) solve for "r"

OpenStudy (jdoe0001):

well. not really bi-weekly in the sense of 52/2 but just twice per month 12 months so 12 * 2 = 24 times a year => "n"

OpenStudy (wolf1728):

Here's a formula for the rate:

OpenStudy (anonymous):

I do not know how to use the formula..

OpenStudy (anonymous):

Math is so hard for me to understand. I plugged the numbers into the formula now im stuck.

OpenStudy (wolf1728):

log(total/principal)/(n*years) I'll work it out for you.

OpenStudy (wolf1728):

log(2,134.96 / 1,327.35) / 24*12 equals log(1.6084378649) / 288 equals 0.2064042885 / 288 log (1 + (rate/n))=0.2064042885 / 288 log (1 + (rate/n))= 0.0007166816

OpenStudy (anonymous):

my choices on my paper are A. 2.9% B. 0.3% C. 4.0% D.4.9%

OpenStudy (wolf1728):

1 + (rate/n) = 10^0.0007166816 1 + (rate/n) = 1.0016515826 Subtracting 1 from both sides rate/n = .0016515826 n = 24 (compounded twice per month) rate = 24*.0016515826 rate = 0.0396379832 rate * 100 = 3.96379832% rounding rate = 4%

OpenStudy (wolf1728):

GlitterBarbie<3 Did you follow that ?

OpenStudy (anonymous):

Kinda, I'm writing it out on paper myself. One second and let me see if i can get it:)

OpenStudy (anonymous):

Was does log mean?

OpenStudy (wolf1728):

It means logarithm. Do you have a log function on your calculator?

OpenStudy (anonymous):

Yes i do.

OpenStudy (wolf1728):

Divide 2,134.96 / 1,327.35 which equals 1.6084378649 then take the logarithm

OpenStudy (wolf1728):

logarithm (1.6084378649) = 0.2064042885

OpenStudy (wolf1728):

I have to get going now. If you want to see another worked out compound interest example, Go to this page http://1728.org/compint3.htm and scroll down to the last problem